OpenStudy (anonymous):

A cone-shaped tent has a diameter of 12 ft and a height of 6 ft. What is the approximate volume of the tent? Use 3.14 to approximate pi, and express your final answer to the nearest tenth. ft3

OpenStudy (anonymous):

@welshfella @e.mccormick

OpenStudy (triciaal):

volume of a cone = 1/3*pi*r^2*h where r is the radius and h is the perpendicular height

OpenStudy (triciaal):

the diameter is twice the radius d = 2r told to use pi as 3.14 to the nearest tenth means round to one decimal place

OpenStudy (anonymous):

@Michele_Laino whats 10851.84 x 1/3 ?

OpenStudy (anonymous):

@triciaal

OpenStudy (michele_laino):

If we use the suggestion of @triciaal we get: \[\Large V = \frac{{\pi \times {6^2} \times 6}}{3} = ...?\]

OpenStudy (michele_laino):

using \pi=3.14, we can write: \[\Large V = \frac{{\pi \times {6^2} \times 6}}{3} = \frac{{3.14 \times {6^2} \times 6}}{3} = ...?\]

OpenStudy (anonymous):

226.08

OpenStudy (michele_laino):

that's right! Now you have to round off that number to the nearest tenth, what do you get?

OpenStudy (anonymous):

226.1

OpenStudy (michele_laino):

that's right!